# Daily-Omni: Towards Audio-Visual Reasoning with Temporal Alignment across Modalities

### v1

@@ 3.2 Data Annotation & QA Construction
We developed a pipeline that employs MLLMs to generate and revise visual and audio annotations. Concurrently, Reasoning Large Language Models (LLMs) are utilized to construct and optimize the associated questions, choices, and answers. To obtain detailed annotations while ensuring cost-effectiveness, we specifically used Gemini 2.0 Flash [noauthor_introducing_2024] for the annotation task and Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] for QA construction and optimization. Figure 3 provides a outline of this process.
@@
Visual & Audio Event Alignment At this stage, we have sequences of visual and audio events annotated within consecutive 10s or 20s segments. However, these segment-level annotations do not explicitly specify the temporal alignment between individual visual and audio events – i.e., which specific events occurred simultaneously. To establish this precise cross-modal event concurrency, we proposed a novel event aligning technique. By prompting Gemini 2.0 Flash with the complete audio-visual clip, we instruct it to identify the visual event(s) occurring concurrently with each identified audio event. These aligned audio-visual event pairs provide sufficient information to infer the temporal relationships between any audio and visual event within the sequence. The details of annotation generation, revision and event alignment is shown in Figure 4.
@@
QA Construction Using the consistent visual annotations, revised audio annotations, and aligned event pairs derived from each video, we prompted Deepseek-R1 to generate multi-choice questions covering the following types: (1) AV Event Alignment: Questions to determine which audio and visual events occurred simultaneously with each other; (2) Event Sequence: Questions to determine the temporal sequence of visual and audio events in the video; (3) Reasoning: Questions to explain the cause or reason behind the occurrence of a visual or audio event in the video; (4) Inference: Questions to speculate on information not explicitly presented in the video; (5) Comparative: Questions to compare the similarity or difference between the audio and visual information of two or more events in the video; (6) Context Understanding: Questions to determine the contextual information surrounding a specific event in the video. Since Daily-Omni aims to evaluate MLLM perception and reasoning within real-life audio-visual scenarios, we deliberately exclude certain question types. Specifically, counting and measuring questions are omitted, as they often rely on a single modality rather than integrated multi-modal understanding. Similarly, purely knowledge-based queries, such as celebrity identification, fall outside our intended scope.
Table 1: Comparison of various audio-visual benchmarks. The table details key characteristics including modality (A: audio, V: video, I: image), dataset size (#Video, #QA Pairs), and Question Type (MC: multiple choice, DF: defined word, BB: bounding boxes). ’Efficient Scalability’ indicates whether the associated work proposed an automated method for benchmark expansion. ’Open-Domain’ denotes benchmarks featuring various categories visual content, while ’General Sound’ signifies benchmarks whose audio content includes speech, music, and generic real-life sounds.
Benchmarks Modality #Video #QA Pairs Question Type Efficient Scalability Open-Domain General Sound
AVQA [avqa] V+A 57,015 57,335 MC
Music-AVQA [li2022learninganswerquestionsdynamic] V+A 9,288 45,867 DF
Pano-AVQA [yun_pano-avqa_2021] V+A 5,400 51,700 DF & BB
OmniBench [li_omnibench_2024] I+A 1,142 MC
AV-Odyssey [gong_av-odyssey_2024] I+A 4,555 MC
WorldSense [hong_worldsense_2025] V+A 1,662 3,172 MC
Daily-Omni (ours) V+A 684 1197 MC
@@
Figure 5: The outline of Daily-Omni Agent.
3.3 Daily-Omni Agent
To establish a baseline for MLLMs and demonstrate the importance of temporal awareness in audio-visual question answering, we constructed the "Daily-Omni Agent". This agent, designed to understand audio-visual context and answer related questions, integrates several models: Qwen2-Audio (7B parameters) [chu2024qwen2audiotechnicalreport], Qwen2.5-VL-7B [bai_qwen25-vl_2025], Whisper-Large-V2 [radford2022robustspeechrecognitionlargescale], and the text-based Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport]. As shown in Figure 5, when presented with a question, its choices, and the associated video context, the agent first divides both the video and audio streams into three segments of equal duration. Subsequently, it independently generates annotations for these segments: visual annotations using Qwen2.5-VL and audio annotations using Qwen2-Audio. Additionally, we use Whisper-Large-V2 to provide a transciption of speech and singing of each segment, as Qwen2-Audio tends to omit this information in its annotations. Following the segment annotation, the agent utilizes Qwen2.5-VL to perform a consistency check on the visual annotations using the complete video, mirroring the revision process described previously (Section 3.2).
While generating aligned audio-visual event pairs would ideally provide richer temporal information for reasoning, implementing this step within the agent presents practical challenges. Unlike the data curation pipeline, the agent lacks access to a highly capable OLM like Gemini 2.0 Flash for precise event alignment. Furthermore, providing a large number of fine-grained event pairs risks overwhelming the context capacity or reasoning capabilities of the agent’s LLM. Therefore, as an alternative to full event alignment, we adopt a innovative targeted approach: First, we prompt Qwen2.5-14B-Instruct with the visual and audio annotations, the question, and its choices. The model’s task is to identify a list of specific visual events whose temporal localization is deemed necessary to answer the question correctly. Subsequently, we utilize Qwen2.5-VL-7B, functioning as a video temporal grounding model, to determine the start and end timestamps for each of these identified critical events. If the duration of a grounded event interval falls below a predefined threshold, the agent classifies this interval as critical. It then retrieves descriptions of both visual and audio events occurring within this specific, brief period using the previously mentioned approach, thereby establishing a localized alignment between concurrent events. Finally, we prompted Qwen2.5-14B-Instruct to determine the correct answer by providing it with the question, choices, the previously generated visual and audio annotations, and the extracted aligned event pairs.
@@ 4 Experiment
In this section, we present a comprehensive evaluation of recent open-source and proprietary Multimodal Large Language Models (MLLMs) on our benchmark to delineate their capability boundaries. Furthermore, we conduct ablation studies to investigate the key factors influencing model performance.
@@ 4.1 Settings
Our evaluation encompassed three distinct types of models. Firstly, we examined OLMs, including several open-source contenders like VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Qwen2.5-Omni [xu_qwen25-omni_2025], Ola [liu_ola_2025] and our Daily-Omni Agent, as well as the proprietary Gemini models [noauthor_introducing_2024]. Secondly, we assessed VLMs, represented by the open-source Qwen2.5-VL [bai_qwen25-vl_2025] and the proprietary GPT-4o [openai_gpt-4o_2024]. We also evaluated the performance of some OLMs when provided with only visual inputs. Finally, we tested some LLMs such as Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [openai_gpt-4o_2024] and Qwen2.5-14B[qwen2025qwen25technicalreport] without visual and audio inputs to check whether our benchmark contains too much information in questions and choices. All model evaluations were conducted strictly according to instructions from their respective official repositories and accompanying documentation (such as "cookbooks" or developer guides).
@@ 4.2 Main Results
Table 2 presents comprehensive evaluation results, shedding light on the real-world audio-visual understanding capabilities of contemporary MLLMs and our proposed Daily-Omni Agent. Firstly, earlier Omni-modal Language Models (OLMs) such as Unified-IO 2 and VideoLLaMA 2 demonstrate limited performance on our benchmarks. Notably, their results are, in several instances, even surpassed by text-only LLMs. Furthermore, the Unified-IO 2 series displays a perplexing degradation in performance with increasing model size, an observation consistent with findings from OmniBench [li_omnibench_2024]. A plausible explanation is an inadequate capacity within these language models to effectively process and synthesize cross-modal information.
Secondly, while overall performance still presents challenges, more recent open-source OLMs such as the Qwen2.5-Omni series and Ola demonstrate reasonable proficiency in inference and reasoning tasks. This suggests a foundational understanding of general visual and audio contexts. However, their efficacy significantly diminishes on temporally-sensitive tasks like audio-visual event alignment and context understanding, where results remain largely unsatisfactory. Despite incorporating cross-modal positional embeddings designed to provide temporal awareness, these models still struggle to address questions requiring precise temporal capabilities. Conversely, proprietary OLMs such as Gemini 2.0 Flash exhibit markedly superior cross-modal temporal capabilities, attaining the leading overall performance score of 67.84%. The observation that even top-tier OLMs have not yet breached the 70% accuracy threshold underscores the demanding nature of this benchmark, positing it as a valuable and pertinent objective for the advancement of contemporary OLMs.
Thirdly, upon ablating audio input and providing only visual data, all evaluated OLMs exhibit a substantial decline in performance. Notably, the magnitude of this performance loss tends to be greater for models that demonstrated superior initial performance when leveraging both modalities. The significant performance drop, especially in high-performing models, when deprived of audio, validates that the tasks within this benchmark genuinely require and benefit from the integration of both auditory and visual information, rather than being solvable by visual cues alone. This dependence on both modalities is further evidenced by the performance of leading Visual Language Models (VLMs). For instance, Qwen2.5-VL (7B) achieves only 40.68% accuracy, and even the highly capable GPT-4o (with its visual input) reaches just 56.47%. Given that these are state-of-the-art models within their respective size categories (sub-7B and larger proprietary models), their comparatively modest performance on our benchmark further underscores its value in demanding genuine audio-visual integration, a capability absent in previous multi-modal benchmarks.
These results reveal that: (1) Current OLMs demonstrate reasonable general audio and visual understanding capabilities when processing videos under 60 seconds in duration. However, they continue to struggle with tasks demanding sophisticated cross-modal temporal awareness or the integration of information across extended time intervals. (2) Our proposed question-answer generation pipeline proves effective in creating challenging evaluation instances that accurately probe the audio-visual understanding capabilities of MLLMs.
Table 2: Performance comparison of various MLLMs across different question domains and video durations. For Omni-modal Language Models (OLMs) with audio and visual inputs, Boldface shows the best performance and underlines show the second best. For OLMs with visual inputs only, subscripts on the ’Avg’ score demonstrate the performance difference compared to their audio-visual counterparts. For your reference, the accuracy rate of pure guessing is 25%. More details of model evaluation are provided in Appendix C.
Methods AV Event Alignment Comparative Context Understanding Event Sequence Inference Reasoning 30s Subset 60s Subset Avg
@@
Qwen2.5-Omni (7B) $44.12$ $51.15$ $38.86$ $40.52$ $57.79$ $61.71$ $46.68$ $48.36$ $47.45$
Qwen2.5-Omni (3B) $38.66$ $48.09$ $33.68$ $33.99$ $54.55$ $44.00$ $42.35$ $38.36$ $40.52$
Ola (7B) $40.34$ $61.07$ $40.41$ $43.46$ $63.64$ $69.71$ $51.47$ $49.82$ $50.71$
Unified-IO-2 L (1B) $27.31$ $22.90$ $26.42$ $27.78$ $29.87$ $29.14$ $27.67$ $27.09$ $27.40$
Unified-IO-2 XL (3B) $30.25$ $30.53$ $25.39$ $29.08$ $33.12$ $21.71$ $28.13$ $28.55$ $28.32$
Unified-IO-2 XXL (8B) $25.63$ $31.30$ $26.42$ $25.82$ $35.06$ $29.71$ $26.74$ $30.00$ $28.24$
VideoLLaMA2 (7B) $35.71$ $35.88$ $35.75$ $31.70$ $40.91$ $34.29$ $38.02$ $31.82$ $35.17$
Gemini 2.0 Flash 62.18 73.28 63.73 63.72 76.62 75.43 67.23 68.55 67.84
Gemini 2.0 Flash Lite 55.04 $64.89$ $58.03$ 54.25 $74.03$ 72.00 $62.44$ 60.00 $61.32$
Daily-Omni (ours) $51.68$ 68.70 60.10 $53.92$ 78.57 $71.43$ 63.99 $59.27$ 61.82
@@
Qwen2.5-Omni (7B) $38.24$ $48.85$ $34.72$ $36.27$ $51.95$ $45.71$ $40.80$ $41.64$ $41.19_{\text{-}6.3}$
Qwen2.5-Omni (3B) $33.61$ $42.75$ $36.27$ $33.33$ $49.35$ $38.86$ $38.79$ $36.55$ $37.76_{\text{-}2.8}$
Gemini 2.0 Flash 39.08 64.12 56.48 56.21 67.53 $62.29$ $56.57$ $55.45$ 56.06-${}_{\text{-}11.8}$
Gemini 2.0 Flash Lite $43.70$ $58.02$ $53.89$ $45.10$ $64.29$ $60.57$ $53.01$ $51.64$ 52.38-${}_{\text{-}8.9}$
@@
GPT-4o $47.90$ $62.60$ $52.33$ $52.61$ 66.23 66.29 55.64 57.45 56.47
Qwen2.5-VL (7B) 36.97 46.56 $33.68$ 37.91 51.95 44.00 39.26 42.36 40.68
Qwen2.5-VL (3B) $35.71$ $43.51$ 34.72 $33.66$ $43.51$ $39.43$ $37.71$ $37.09$ $37.43$
@@
GPT-4o $33.19$ $43.51$ $28.50$ $30.39$ $44.81$ $46.86$ $36.48$ $36.18$ $36.34$
Deepseek-V3 (671B) $31.93$ $41.22$ $29.02$ $29.41$ $44.81$ $46.29$ $35.24$ $36.00$ $35.59$
Qwen2.5-Instruct (14B) $30.25$ $39.69$ $27.98$ $28.43$ $42.21$ $42.86$ $32.15$ $35.82$ $33.83$
4.3 Results of Daily-Omni Agent
Our Daily-Omni Agent, leveraging Qwen2.5-VL-7B, Qwen2-Audio (7B), and Qwen2.5-Instruct-14B, achieves 61.82% overall accuracy. This performance is state-of-the-art among open-source methods and surpasses smaller proprietary MLLMs. The success of our agent suggests that by simply leveraging VLM and ALM with time-period-level alignment coupled with event-level alignment for some key events, we are able to create a very strong OLM.

### v2

@@ 3.2 Data Annotation & QA Construction
We developed a pipeline that employs MLLMs to generate and revise visual and audio annotations. Concurrently, Reasoning Large Language Models (LLMs) are utilized to construct and optimize the associated questions, choices, and answers. To obtain detailed annotations while ensuring cost-effectiveness, we specifically used Gemini 2.0 Flash [noauthor_introducing_2024] for the annotation task and Deepseek-R1 [deepseekai2025deepseekr1incentivizingreasoningcapability] for QA construction and optimization. Figure 3 provides an outline of this process.
@@
Visual & Audio Event Alignment At this stage, we have sequences of visual and audio events annotated within consecutive 10s or 20s segments. However, these segment-level annotations do not explicitly specify the temporal alignment between individual visual and audio events, i.e., which specific events occurred simultaneously. To establish this precise cross-modal event concurrency, we proposed a event aligning technique. By prompting Gemini 2.0 Flash with the complete audio-visual clip, we instruct it to identify the visual event(s) occurring concurrently with each identified audio event. These aligned audio-visual event pairs provide sufficient information to infer the temporal relationships between any audio and visual event within the sequence. The details of annotation generation, revision and event alignment is shown in Figure 4. We further assess the reliability of this alignment step via manual verification on a randomly sampled subset of 100 videos: over 90% of aligned audio–visual event pairs are judged correct.
@@
QA Construction Using the consistent visual annotations, revised audio annotations, and aligned event pairs derived from each video, we prompted Deepseek-R1 to generate multi-choice questions covering the following types: (1) AV Event Alignment: To determine which audio and visual events occurred simultaneously with each other; (2) Event Sequence: To determine the temporal sequence of visual and audio events in the video; (3) Reasoning: To explain the cause or reason behind the occurrence of a visual or audio event in the video; (4) Inference: To speculate on information not explicitly presented in the video; (5) Comparative: To compare the similarity or difference between the audio and visual information of two or more events in the video; (6) Context Understanding: To determine the contextual information surrounding a specific event in the video. Since Daily-Omni aims to evaluate MLLM perception and reasoning within real-life audio-visual scenarios, we deliberately exclude certain question types. Specifically, counting and measuring questions are omitted, as they often rely on a single modality rather than integrated multi-modal understanding. Similarly, purely knowledge-based queries such as celebrity identification are not included.
@@
Figure 5: The outline of Daily-Omni Agent workflow.
3.3 Diagnostic Baseline: Daily-Omni Agent
To establish a heuristic reference for MLLMs and diagnose the importance of explicit temporal awareness in audio-visual question answering, we constructed a training-free diagnostic pipeline, termed the "Daily-Omni Agent". This baseline, designed to understand audio-visual context and answer related questions, integrates several models: Qwen2-Audio [chu2024qwen2audiotechnicalreport], Qwen2.5-VL-7B [bai_qwen25-vl_2025], Whisper-Large-V2 [radford2022robustspeechrecognitionlargescale], and the text-based Qwen2.5-14B-Instruct [qwen2025qwen25technicalreport]. As shown in Figure 5, when presented with a question, its choices, and the associated video context, the agent first divides both the video and audio streams into three segments of equal duration. Subsequently, it independently generates annotations for these segments: visual annotations using Qwen2.5-VL and audio annotations using Qwen2-Audio. Additionally, we use Whisper-Large-V2 to provide a transciption of speech and singing of each segment, as Qwen2-Audio tends to omit this in its annotations. Following the segment annotation, the agent utilizes Qwen2.5-VL to perform a consistency check on the visual annotations using the complete video, mirroring the revision process described previously (Section 3.2).
While generating aligned audio-visual event pairs would ideally provide richer temporal information for reasoning, implementing this step within the agent presents practical challenges. Unlike the data curation pipeline, the agent lacks access to a highly capable OLM like Gemini 2.5 Flash for precise event alignment. Furthermore, providing a large number of fine-grained event pairs may overwhelm the context capacity or reasoning capabilities of the agent’s LLM. Therefore, we adopt a targeted approach: First, we prompt Qwen2.5-14B-Instruct with the visual and audio annotations, the question, and its choices. The model’s task is to identify a list of specific visual events whose temporal localization is deemed necessary to answer the question correctly. Subsequently, we utilize Qwen2.5-VL-7B, functioning as a video temporal grounding model, to determine the start and end timestamps for each of these identified critical events. If the duration of a grounded event interval falls below a predefined threshold, the agent classifies this interval as critical. It then retrieves descriptions of both visual and audio events occurring within this specific, brief period using the previously mentioned approach, thereby establishing a localized alignment between concurrent events. Finally, we prompted Qwen2.5-14B-Instruct to determine the correct answer by providing it with the question, choices, the previously generated visual and audio annotations, and the extracted aligned event pairs.
@@ 4 Experiment
In this section, we present a comprehensive evaluation of recent Multimodal Large Language Models (MLLMs) on our benchmark to delineate their capability boundaries. Furthermore, we conduct ablation studies to investigate the key factors influencing model performance.
@@ 4.1 Settings
Our evaluation encompassed three distinct types of models. Firstly, we examine OLMs, including several open-source contenders like VideoLLaMA 2 [cheng_videollama_2024], Unified-IO 2 [lu2023unifiedio2scalingautoregressive], Qwen2.5-Omni [xu_qwen25-omni_2025], Qwen3-Omni [Qwen3-Omni], Ola [liu_ola_2025] and our Daily-Omni Agent, as well as the proprietary Gemini 2.0 and 2.5 models [noauthor_introducing_2024, gemini25]. Secondly, we assess VLMs including Qwen2.5-VL [bai_qwen25-vl_2025] and GPT-4o [openai_gpt-4o_2024] and ALMs including Audio Flamingo 3 [goel2025audioflamingo3advancing] and Qwen2-Audio [chu2024qwen2audiotechnicalreport]. We also evaluate the performance of some OLMs when provided with only visual inputs. Finally, we test some LLMs such as Deepseek-V3 [deepseekai2025deepseekv3technicalreport], GPT-4o [openai_gpt-4o_2024] and Qwen2.5-14B[qwen2025qwen25technicalreport] without visual and audio inputs to evaluate the extent of textual bias in our questions and choices. All model evaluations are conducted strictly according to instructions from their respective official repositories and accompanying documentation (such as "cookbooks" or developer guides).
@@ 4.2 Main Results
Table 2: Performance comparison of MLLMs on Daily-Omni. Boldface and underline indicate the top two performers. Subscripts on ’Avg’ for visual-only OLMs show the performance drop from their audio-visual counterparts. Abbreviations: AV Align = audio-visual alignment, Comp. = comparative, Ctx. Und. = context understanding, Evt. Seq. = event sequence, Infer. = inference, Reas. = reasoning, and 30s/60s denote duration-based subsets. Random guess accuracy is 25%. Color code: Gray background indicates closed-source models and blue background indicates open-source models.
Methods AV Align Comp. Ctx. Und. Evt. Seq. Infer. Reas. 30s 60s Avg
@@
[…]
@@
[…]
Omni-Modal Language Models (Audio Only)
[…]
@@
[…]
Audio Language Models (Audio Only)
[…]
@@
[…]
Table 2 presents comprehensive evaluation results, shedding light on the real-world audio-visual understanding capabilities of contemporary MLLMs and our diagnostic reference baseline, the Daily-Omni Agent. Firstly, earlier Omni-modal Language Models (OLMs) such as Unified-IO 2 and VideoLLaMA 2 demonstrate limited performance on our benchmark. Notably, their results are, in several instances, even surpassed by text-only LLMs. Furthermore, the Unified-IO 2 series displays a perplexing degradation in performance with increasing model size, an observation consistent with findings from OmniBench [li_omnibench_2024]. A plausible explanation is that these models may not reliably synthesize cross-modal information under our evaluation setting.
Secondly, we observe a clear performance divide among recent OLMs, highlighting the importance of cross-modal temporal alignment for audio-visual QA. Our training-free Daily-Omni Agent, which makes temporally localized evidence more explicit via a decoupled pipeline, serves as a useful reference point with an overall accuracy of 61.82%. Models that perform substantially below this reference (such as Ola and earlier OLMs) appear to struggle with synchronizing parallel audio and visual streams for alignment-sensitive questions. In contrast, models that achieve higher performance demonstrate stronger proficiency in audio-visual reasoning. While proprietary models like Gemini 2.5 Flash achieve top-tier performance (up to 73.06%), among open-source models, the Qwen2.5-Omni and Qwen3-Omni series stand out. Their gains likely reflect a combination of factors, including training data and objectives, architectural choices for multimodal fusion, and temporal modeling components. In particular, explicit cross-modal positional designs (e.g., TMRoPE) may be one contributing technique for improving temporal awareness across modalities. Overall, these results suggest that robust spatio-temporal alignment mechanisms remain an important ingredient for tackling complex audio-visual reasoning tasks.
These results reveal that: (1) Advanced OLMs with stronger temporal handling (e.g., the Qwen Omni series) or proprietary systems (e.g., Gemini) demonstrate solid audio-visual understanding, though they still face challenges in complex temporal reasoning. (2) Models with weaker temporal integration tend to perform poorly on alignment-sensitive questions, even when provided with both audio and visual inputs. (3) Our QA construction and filtering pipeline produces challenging evaluation instances that probe the temporal reasoning boundaries of current MLLMs.
Figure 6: Evaluation stability under video-level subsampling. Lines show median overall accuracy and shaded bands denote the 5th–95th percentile across 200 trials.
4.3 Evaluation Stability via Subsampling
To assess the stability of aggregate evaluation on Daily-Omni, we conduct a video-level subsampling study on representative OLMs. We repeatedly sample {20,40,60,80}% of videos without replacement and evaluate overall accuracy on the induced QA subsets (Figure 6). The estimates remain consistent across subsamples and the variability decreases rapidly as more videos are included: at 80% of videos (approximately 958 questions), the 5th–95th percentile range is within 1.1–1.2 percentage points for Gemini 2.5 Flash, Gemini 2.5 Flash Lite, and Qwen3-Omni-30B. Overall, this subsampling analysis supports that Daily-Omni provides sufficiently stable aggregate evaluation to serve as a robust benchmark.
4.4 Ablation Study
To further validate the design of our proposed baseline and the intrinsic multimodal requirements of the Daily-Omni benchmark, we conduct comprehensive ablation studies focusing on temporal alignment methods and modality dependency.
(a) Modality ablation on Daily-Omni
(b) Impact of alignment methods
Figure 7: Ablation studies validating the design of the benchmark and the baseline agent. (a) Performance comparison of representative OLMs under full modalities (A+V), Audio-only, and Visual-only settings. (b) Comparison of average accuracy (bars) and aligned event pairs per question (line) using our baseline agent.
4.4.1 Modality Ablation
To verify whether the tasks in Daily-Omni genuinely require the integration of both auditory and visual information, we evaluated representative top-performing OLMs (Qwen3-Omni-30B-A3B-Instruct/Thinking, Gemini 2.5 Flash, and Gemini 2.5 Flash Lite) under three conditions: full modalities (Audio + Visual), audio-only, and visual-only. The comparative results are illustrated in Figure 7(a).
Upon ablating either modality, all evaluated OLMs exhibit a substantial decline in performance. For instance, the highly capable Gemini 2.5 Flash experiences a drastic accuracy drop from 73.06% to 54.05% without visual input, and down to 44.61% without audio input. Similarly, the Qwen3-Omni models suffer a 13–16% performance degradation when deprived of either modality. This significant performance gap between the full-modality and uni-modal settings strongly validates that the questions in our benchmark cannot be reliably answered using a single sensory channel or text-based guessing.
Interestingly, we observe that for these models, the audio-only performance is generally higher than the visual-only performance on our dataset. This observation highlights that unlike many existing video QA benchmarks that exhibit a strong visual bias (where audio is often redundant), Daily-Omni contains rich, indispensable acoustic events and spoken dialogues. In conclusion, this ablation confirms that achieving high performance on Daily-Omni strictly necessitates genuine multi-modal synergy and precise cross-modal integration.
4.4.2 Daily-Omni Agent Alignment
Our heuristic Daily-Omni Agent, leveraging Qwen2.5-VL-7B, Qwen2-Audio (7B), and Qwen2.5-Instruct-14B, achieves 61.82% overall accuracy. The fact that this simple, decoupled pipeline surpasses smaller proprietary MLLMs and several unified open-source models strongly validates the core motivation of our benchmark: temporal alignment is the primary bottleneck in audio-visual reasoning.
To further study the effect of event alignment, we conducted an ablation study evaluating the Daily-Omni Agent’s performance under three alignment scenarios: (1) No Alignment: Generating comprehensive visual and audio captions as a base, but omitting any form of event alignment. (2) Naive Alignment: Generating comprehensive visual and audio captions, then aligning events in a question-agnostic manner. Specifically, we prompt the VLM to enumerate visual events from the captions and localize each event by predicting its temporal boundaries (start/end). For each localized interval, we retrieve the corresponding audio events within the same time range using the ALM, and form aligned audio–visual event pairs. (3) Smart Alignment: Generating comprehensive visual and audio captions and aligned events as stated in Section 3.3. Note that for (2) and (3), we only consider generating an align event pair when the provided segment has a duration below a certain threshold to prevent confusion.